OpenStudy (anonymous):

follow principles of estimation to find an approximate value. round each number so that there is one nonzero digit do not find the exact value 12 Division 493,867 I believe you round down the 12 to 10 is the 493,867 rounded up or down please thanks in advance

OpenStudy (phi):

they want only one non-zero digit, which means you will change 493,867 to the either 400,000 or 500,000 you should pick the one that is nearest to 493,867

OpenStudy (anonymous):

so with the 3rd digit below 5 it should be brought down to 400,000

OpenStudy (phi):

in general, you would look at the digit to the right of the 4... in this case the 9 if that digit is 5 or more, round up. 9 is bigger than 5, so round 4 up to 5 the rounded number is 500,000

OpenStudy (anonymous):

ok that makes sense

OpenStudy (phi):

you could write 493,867 as 400,000 + 93,867 notice that 93,867 is getting close to 100,000 (you are only 6,133 away) so this is close to 400,000 + 100,000 = 500,000

OpenStudy (phi):

***so with the 3rd digit below 5 it should be brought down to 400,000*** you look at the 2nd digit in, not the 3rd digit.

OpenStudy (anonymous):

great explanation thank you

OpenStudy (phi):

yw

OpenStudy (phi):

is the problem to estimate 12/493,867 ?

OpenStudy (phi):

the reason I asked is because there is a difference between 10/500,000 = 1/50,000 (a tiny number) and 500,000 / 10 = 50,000 (a big number)